OpenStudy (mtepecz):

Use the given information to find exact answers for cos(x/2), sin(x/2), and tan(x/2): Cos (x)=-4/5 180 degrees< x < 270 degrees

OpenStudy (anonymous):

\[2 \sin ^2 \frac{ x }{ 2 }=1-\cos x\] \[2 \cos ^2\frac{ x }{ 2 }=1+\cos x\] \[180<x<270,90<\frac{ x }{ 2 }<135\] \[\frac{ x }{ 2 }~lies~\in~second~quadrant.\] \[\sin \frac{ x }{ 2 }~is~positive.\] \[\cos \frac{ x }{ 2 }~is~negative.\] \[\tan \frac{ x }{ 2 }~is~negative.\] \[\tan \frac{ x }{ 2 }=\frac{ \sin \frac{ x }{ 2 } }{ \cos \frac{ x }{ 2 } }\]